# Gappy Data Reconstruction using Unsupervised Learning for Digital Twin

### 13 Dec 2023 v1

@@ 2.3 Sampling Algorithms
(2.18) $\mathbf{r}=\mathbf{u}-\boldsymbol{g}(\tilde{\mathbf{u}}),$
where $\boldsymbol{g}:\mathbb{R}^{{n_{s}}}\rightarrow\mathbb{R}^{N_{s}}$ is a nonlinear function, assuming $\mathbf{u}_{ref}=0$. Following the GNAT-SNS approach introduced in [15], the residual is approximated by
@@
(2.24) $\hat{\mathbf{u}}=\underset{\hat{\boldsymbol{v}}\in\mathbb{R}^{{n_{s}}}}{\text{argmin}}\lVert\boldsymbol{\Phi}_{r}(\boldsymbol{Z}^{T}\boldsymbol{\Phi}_{r})^{-1}\boldsymbol{Z}^{T}(\mathbf{u}-\boldsymbol{g}(\hat{\boldsymbol{v}}))\rVert_{2}^{2}$
solves $\hat{\mathbf{u}}$ using the Gauss–Newton method. The update for each iteration is given by
(2.25) $\hat{\mathbf{u}}_{i}=\hat{\mathbf{u}}_{i-1}+\{(\boldsymbol{Z}^{T}\boldsymbol{\Phi}_{r})^{-1}\boldsymbol{Z}^{T}\mathbf{J}(\hat{\mathbf{u}}_{i-1})\}^{\dagger}(\boldsymbol{Z}^{T}\boldsymbol{\Phi}_{r})^{-1}\boldsymbol{Z}^{T}(\mathbf{u}-\boldsymbol{g}(\hat{\mathbf{u}}_{i-1})),$
where $\mathbf{J}(\hat{\mathbf{u}})$ is a Jacobian of $\boldsymbol{g}(\hat{\mathbf{u}})$ and ${}^{\dagger}$ denotes the Moore–Penrose inverse of a matrix.
@@ 2.3.2 S-OPT
We introduce the S-OPT sampling algorithm presented in [46, 35]. As in Section 2.3.1, we denote $\boldsymbol{\Phi}$ as the POD basis matrix, $\boldsymbol{Z}^{T}$ as the sampling matrix, and $\boldsymbol{u}$ as the ground truth data to be reconstructed. According to Theorem 3.1 of [35], we can write the reconstruction error as
(2.28) $\|(\boldsymbol{I}-\boldsymbol{\Phi}(\boldsymbol{Z}^{T}\boldsymbol{\Phi})^{\dagger}\boldsymbol{Z}^{T})\boldsymbol{u}\|_{2}^{2}=\|(\boldsymbol{I}-\boldsymbol{\Phi}\boldsymbol{\Phi}^{T})\boldsymbol{u}\|_{2}^{2}+\|\boldsymbol{\epsilon}\|_{2}^{2},$
where $\boldsymbol{\epsilon}=((\boldsymbol{Z}^{T}\boldsymbol{\Phi})^{T}\boldsymbol{Z}^{T}\boldsymbol{\Phi})^{-1}(\boldsymbol{Z}^{T}\boldsymbol{\Phi})^{T}\boldsymbol{Z}^{T}(\boldsymbol{I}-\boldsymbol{\Phi}\boldsymbol{\Phi}^{T})\boldsymbol{u}$. Since we do not know the ground truth of $\boldsymbol{u}$ in practice, the true optimum $\boldsymbol{Z}^{\ast}$ cannot be found. We use training data to find $\boldsymbol{Z}$ and this is quasi-optimal. The S-OPT algorithm finds sampling points to have the smallest $\boldsymbol{\epsilon}$ under the condition of given training data of $\boldsymbol{u}$, corresponding to maximizing the column orthogonality of $\boldsymbol{Z}^{T}\boldsymbol{\Phi}$ and determinant of $(\boldsymbol{Z}^{T}\boldsymbol{\Phi})^{T}\boldsymbol{Z}^{T}\boldsymbol{\Phi}$. In [46], the quantity $S:\mathbb{R}^{N\times p}\rightarrow[0,1]$ is defined as
@@ 3 Numerical Results
The mean squared error (MSE) loss of the auto-encoder is shown in Fig. 1 of Appendix A in [31]. The training loss is lower than the testing loss for all cases. We conclude that the auto-encoder models are not over-fitted to the training data.
Considering randomness due to random initialization and mini-batch sampling, The auto-encoder training is repeated 10 times to account for randomness due to random initialization and mini-batch sampling. The statistics of MSE losses are presented in Fig. 2 of Appendix A in [31]. The MSE losses decrease as latent space dimensions increase because of more learnable hyper-parameters.
@@
For our numerical examples, the size of the reduced dimension (i.e., latent space dimension for nonlinear manifold or the reduced basis dimension for linear subspace) should be at most half of the number of sample points to avoid numerical round-off errors. In addition, the reduced dimension size is greater than that of the intrinsic dimension. Through the numerical examples, the intrinsic dimension is 2, with time $t\in\mathbb{R}_{+}$ and parameter $\mu\in\mathbb{R}$; the number of sample points is 12. Thus, we limit the dimension size to $\{3,4,5,6\}$.
The computational cost of the online phase of gappy data reconstruction is measured in terms of CPU wall time. Calculations are performed on an AMD EPYC 7763 CPU @ 2.45 GHz and DDR4 Memory @ 3200 MT/s. With given $12$ sample points and reduced dimension (i.e., the POD basis or latent space dimensions) of $[3,4,5,6]$, the online phase of the Gappy POD method for each reconstruction step in our example problems takes around $1$ to $2$ ms. However, it takes around $15$ to $18$ ms for the Gappy AE method to reconstruct the data during the online phase. The performances of Gappy AE and POD methods in terms of reconstruction error are compared in the following sections.
@@ 3.1 Diffusion Problem
where spatial and temporal domains are given by $\boldsymbol{x}\in\Omega=[0,1]\times[0,1]$ and $t\in[0,1]$, respectively. Here, we set $\kappa=0.05$ and $\alpha=0.01$. The Neumann boundary condition is imposed
(3.4) $\frac{\partial u}{\partial\boldsymbol{x}}\cdot\boldsymbol{n}=0\quad\text{on}\quad\partial\Omega,$
@@ 3.2 Radial Advection Problem
where the spatial and time domains are given by $\boldsymbol{x}\in\Omega=[-1,1]\times[-1,1]$ and $t\in[0,3]$, respectively; velocity, $\mathbf{v}$ is defined as
(3.7) $\mathbf{v}=\frac{\pi}{2}(1-x_{1}^{2})^{2}(1-x_{2}^{2})^{2}(x_{2},-x_{1})^{T}.$
@@ 4 Discussion & conclusion
We proposed the Gappy AE method for gappy data reconstruction and compared it with the popular Gappy POD method. In the Gappy AE method, data is reconstructed using a nonlinear manifold by training the auto-encoder. Unlike a linear subspace-based approach, such as the Gappy POD method, the Gappy AE shows $\mathcal{O}(10)$ to $\mathcal{O}(100)$ times less data reconstruction error when assuming noiseless measurement. The superior representability of the nonlinear manifold allows the Gappy AE to reconstruct gappy data when known (measured) data is sparse. This prominent feature is aligned with the digital twin application for physical data because, in reality, sensors are limited to measuring sparse physical field data, yet they want to estimate the entire field data. In our numerical tests, the spatial dimension of reconstruction data is $3,600$ and the number of sample points is $12$. The speed of data reconstruction using the Gappy POD is approximately $10$ times faster than the Gappy AE. Although the Gappy AE takes more time to reconstruct data because of solving nonlinear problems in latent space dimension, its speed is sufficient for a $60$ Hz measurement frequency.
@@
In future work, we will investigate retraining the auto-encoder with noisy data to make the Gappy AE more reliable for noisy measurements.

### 31 Mar 2024 v2

@@ 2.3 Sampling Algorithms
(2.19) $\mathbf{r}=\boldsymbol{x}-\boldsymbol{g}(\hat{\boldsymbol{x}}),$
assuming $\boldsymbol{x}_{ref}=0$. Following the GNAT-SNS approach introduced in [15], the residual is approximated by
@@
(2.25) $\hat{\boldsymbol{x}}=\underset{\hat{\boldsymbol{v}}\in\mathbb{R}^{{n_{s}}}}{\text{argmin}}\lVert\boldsymbol{\Phi}_{r}(\boldsymbol{Z}^{T}\boldsymbol{\Phi}_{r})^{-1}\boldsymbol{Z}^{T}(\boldsymbol{x}-\boldsymbol{g}(\hat{\boldsymbol{v}}))\rVert_{2}^{2}$
solves $\hat{\boldsymbol{x}}$ using the Gauss–Newton method. The update for each iteration is given by
(2.26) $\hat{\boldsymbol{x}}_{i}=\hat{\boldsymbol{x}}_{i-1}+\{(\boldsymbol{Z}^{T}\boldsymbol{\Phi}_{r})^{-1}\boldsymbol{Z}^{T}\mathbf{J}(\hat{\boldsymbol{x}}_{i-1})\}^{\dagger}(\boldsymbol{Z}^{T}\boldsymbol{\Phi}_{r})^{-1}\boldsymbol{Z}^{T}(\boldsymbol{x}-\boldsymbol{g}(\hat{\boldsymbol{x}}_{i-1})),$
where $\mathbf{J}(\hat{\boldsymbol{x}})$ is a Jacobian of $\boldsymbol{g}(\hat{\boldsymbol{x}})$.
@@ 2.3.2 S-OPT
We introduce the S-OPT sampling algorithm presented in [48, 36]. As in Section 2.3.1, we denote $\boldsymbol{\Phi}$ as the POD basis matrix, $\boldsymbol{Z}^{T}$ as the sampling matrix, and $\boldsymbol{x}$ as the ground truth data to be reconstructed. According to Theorem 3.1 of [36], we can write the reconstruction error as
(2.29) $\|(\boldsymbol{I}-\boldsymbol{\Phi}(\boldsymbol{Z}^{T}\boldsymbol{\Phi})^{\dagger}\boldsymbol{Z}^{T})\boldsymbol{x}\|_{2}^{2}=\|(\boldsymbol{I}-\boldsymbol{\Phi}\boldsymbol{\Phi}^{T})\boldsymbol{x}\|_{2}^{2}+\|\boldsymbol{\epsilon}\|_{2}^{2},$
where $\boldsymbol{\epsilon}=((\boldsymbol{Z}^{T}\boldsymbol{\Phi})^{T}\boldsymbol{Z}^{T}\boldsymbol{\Phi})^{-1}(\boldsymbol{Z}^{T}\boldsymbol{\Phi})^{T}\boldsymbol{Z}^{T}(\boldsymbol{I}-\boldsymbol{\Phi}\boldsymbol{\Phi}^{T})\boldsymbol{x}$. Since we do not know the ground truth of $\boldsymbol{x}$ in practice, the true optimum $\boldsymbol{Z}^{\ast}$ cannot be found. We use training data to find $\boldsymbol{Z}$ and this is quasi-optimal. The S-OPT algorithm finds sampling points to have the smallest $\boldsymbol{\epsilon}$ under the condition of given training data of $\boldsymbol{x}$, corresponding to maximizing the column orthogonality of $\boldsymbol{Z}^{T}\boldsymbol{\Phi}$ and determinant of $(\boldsymbol{Z}^{T}\boldsymbol{\Phi})^{T}\boldsymbol{Z}^{T}\boldsymbol{\Phi}$. In [48], the quantity $S:\mathbb{R}^{N\times p}\rightarrow[0,1]$ is defined as
@@ 3 Numerical Results
In our numerical examples, the width of the encoder is twice the input size and the latent space dimension varies from 3 to 6. The block size and the amount of shifts are set as 100 and $20$, respectively. The batch size is 240 and the data set is split into $80\%$ for training and $20\%$ for testing. The ADAM optimizer, a variant of stochastic gradient descent (SGD), is used to update the learnable parameters. The initial learning rate is set as 0.001 and reduced by a factor of 10 if learning shows no improvement for 50 epochs. The number of epochs for training is 10000, but if training does not yield improvement after 200 times, the training is stopped. The autoencoders are trained on a single NVIDIA A100 GPU with 40GB of memory. The average training time per epoch is $0.539$ seconds with a latent space dimension set to $6$. The early stopping scheme introduces variability in the training time; however, we enforce a maximum of $10000$ epochs. Consequently, the training time is capped at approximately $90$ minutes. The mean squared error (MSE) loss of the auto-encoder is shown in Fig. 1 of Appendix A in [32]. For all cases, both training curves and test curves consistently continue to improve and we don’t see the test curves plateau or worsen, which suggests the models are not overfitting. We conclude that the auto-encoder models are not over-fitted to the training data.
The auto-encoder training is repeated 10 times to account for randomness due to random initialization and mini-batch sampling. The statistics of MSE losses are presented in Fig. 2 of Appendix A in [32]. The MSE losses decrease as latent space dimensions increase because of more learnable hyper-parameters.
@@
For our numerical examples, the size of the reduced dimension (i.e., latent space dimension for nonlinear manifold or the reduced basis dimension for linear subspace) should be at most half of the number of sample points to avoid numerical round-off errors. If the reduced dimension size is close to the number of sample points, the condition number of the matrix in solving $\hat{\boldsymbol{x}}$ increases, leading to errors in the calculation of the pseudo-inverse. In addition, the reduced dimension size is greater than that of the intrinsic dimension. Through the numerical examples, the intrinsic dimension is 2, with time $t\in\mathbb{R}_{+}$ and parameter $\mu\in\mathbb{R}$; the number of sample points is 12. Thus, we limit the dimension size to $\{3,4,5,6\}$.
The computational cost of the online phase of gappy data reconstruction is measured in terms of wall-clock time. Calculations are performed on one CPU of an AMD EPYC 7763 @ 2.45 GHz and DDR4 Memory @ 3200 MT/s. With given $12$ sample points and reduced dimension (i.e., the POD basis or latent space dimensions) of $[3,4,5,6]$, the online phase of the Gappy POD method for each reconstruction step in our example problems takes around $0.1$ ms. However, it takes around $3.0$ to $3.8$ ms for the Gappy AE method to reconstruct the data during the online phase. To compare the online cost of Gappy AE, Gappy POD, and FE simulation, wall-clock time for each time step is measured $10$ times and the mean value is used in Table 1. The performances of Gappy AE and POD methods in terms of reconstruction error are compared in the following sections.
Table 1: Comparison of online cost of Gappy AE and Gappy POD under noiseless measurement conditions. Note that FE simulation does not reconstruct data but solves governing equations. For data reconstruction, the best-performing cases in terms of accuracy from each numerical example are selected, and the online costs of Gappy AE, Gappy POD, and FE simulation are computed by averaging the mean wall-clock times for each time step over FOM parameters.
Problem Type Sensor Placement Time (ms)
Gappy AE Gappy POD FE simulation
Diffusion Boundary 3.794 0.119 44.481
Domain 3.045 0.118
Radial Advection Domain 3.161 0.121 32.220
Wave Boundary 3.557 0.121 59.400
Domain 3.355 0.121
@@ 3.1 Diffusion Problem
where spatial and temporal domains are given by $\mathbf{x}\in\Omega=[0,1]\times[0,1]$ and $t\in[0,1]$, respectively. Here, we set $\kappa=0.05$ and $\alpha=0.01$. The Neumann boundary condition is imposed
(3.4) $\frac{\partial u}{\partial\mathbf{x}}\cdot\boldsymbol{n}=0\quad\text{on}\quad\partial\Omega,$
@@ 3.2 Radial Advection Problem
where the spatial and time domains are given by $\mathbf{x}\in\Omega=[-1,1]\times[-1,1]$ and $t\in[0,3]$, respectively; velocity, $\mathbf{v}$ is defined as
(3.7) $\mathbf{v}=\frac{\pi}{2}(1-\mathrm{x}_{1}^{2})^{2}(1-\mathrm{x}_{2}^{2})^{2}(\mathrm{x}_{2},-\mathrm{x}_{1})^{T}.$
@@ 4 Discussion & conclusion
We proposed the Gappy AE method for gappy data reconstruction and compared it with the popular Gappy POD method. In the Gappy AE method, data is reconstructed using a nonlinear manifold by training the auto-encoder. Unlike a linear subspace-based approach, such as the Gappy POD method, the Gappy AE shows $\mathcal{O}(10)$ to $\mathcal{O}(100)$ times less data reconstruction error when assuming noiseless measurement as shown in Table 2.
Table 2: Comparison of relative errors between Gappy AE and Gappy POD under noiseless measurement conditions. The best-performing cases from each numerical example are selected, and their relative errors are computed by averaging over FOM parameters.
Problem Type Sensor Placement Relative Error (%)
Gappy AE Gappy POD
Diffusion Boundary 0.0896 9.7730
Domain 0.0753 2.6964
Radial Advection Domain 0.0699 7.8696
Wave Boundary 0.4102 2.5992
Domain 0.3272 1.6759
Even for diffusion dominant problems, the Gappy AE can outperform Gappy POD in terms of accuracy when the number of sample points is very sparse. However, we acknowledge that Gappy POD will in general outperform Gappy AE in terms of speed-up. The superior representability of the nonlinear manifold allows the Gappy AE to reconstruct gappy data when known (measured) data is sparse. This prominent feature is aligned with the digital twin application for physical data because, in reality, sensors are limited to measuring sparse physical field data, yet they want to estimate the entire field data. In addition to that, if there is training data available for autoencoders, the Gappy AE method can be used in any domains whether simulation is possible or not. In our numerical tests, the spatial dimension of reconstruction data is $3,600$ and the number of sample points is $12$. The speed of data reconstruction using the Gappy POD is approximately $30$ times faster than the Gappy AE. Although the Gappy AE takes more time to reconstruct data because of solving nonlinear problems in latent space dimension, its speed is sufficient for a $260$ Hz measurement frequency.
@@
In future work, we will investigate retraining the auto-encoder with noisy data to make the Gappy AE more reliable for noisy measurements. To be more specific, we plan to train two decoders. One is for noisy data and the other is for noiseless data. The decoder trained with noisy data can be used for finding $\hat{\boldsymbol{x}}$ in a latent space. Then, the decoder trained with noiseless data will be used to reconstruct data.